I need help figuring out how to solve this.
It is an infinite series, 1+3+9+27+81+243+...= -1/2
I need to prove that this infinite series equals -1/2.

Answer by Alan3354(69443)

I need help figuring out how to solve this.
It is an infinite series, 1+3+9+27+81+243+...= -1/2
I need to prove that this infinite series equals -1/2.
-----------
It obviously does not = -1/2.
Why would it be negative?
The 1st term is > -1/2